# Chen Xu1,2, Zhenyu Lv1,2, Tian Lan4, Xianyang Wang1,2, Luyao Ji3, Leyang Cui Minqiang Yang5, Jian Shen1,2, Qunxi Dong1,2, Xiuling Liu6, Juan Wang3, Bin Hu1,2,5

### v1

@@ Introduction
The field of mental health is grappling with a profound imbalance between the increasing number of patients and the scarcity of available therapists (WHO 2023). Although Large language models (LLMs) offer a potential solution to this imbalance by simulating human therapist behaviors and providing personalized counseling services at lower costs (Xiao et al. 2024; Na 2024; Xie et al. 2024; Wang et al. 2025) , their direct application in patient-facing treatment poses several risks. These risks include ethical dilemmas (Ong et al. 2024), safety concerns (Hager et al. 2024), potential harm with uncertainty about responsibility (Nichol et al. 2023), and the inherent challenge of establishing a therapeutic alliance with virtual therapists (Choudhury and Chaudhry 2024). This prompts us to explore automated methods for training or supervising therapists to accelerate their development, addressing the imbalance by empowering more therapists through AI-assisted means.
@@
With significant advances in LLM role play  (Xu et al. 2022; Tseng et al. 2024; Chen et al. 2024), a new AI-assisted training paradigm has emerged (Yang et al. 2024). In this paradigm, LLMs play partners in the field and collaborate with practitioners to create interactive scenarios for experiential learning (Kolb 2014). For example, in educational scenarios, “AI Student as Partner” helps teaching assistants improve their communication and classroom management skills (Markel et al. 2023); in conflict resolution, “AI Opponent as Partner” helps users practice negotiation and conflict resolution (Shaikh et al. 2023). However, in psychotherapy, “AI Patient as Partner” (Zhang et al. 2025) is difficult to implement. Privacy concerns and limited clinical data hinder the accurate modeling of patients’ behavioral patterns, as well as the already intricate cognitive states (Reger et al. 2021; Lee et al. 2024a).
@@
Recently, a pioneering work, Patient-$\Psi$-Trainer has been proposed to fill this gap (Wang et al. 2024). It integrates 106 cognitive models of patients (Beck 2020), manually created by clinical experts, into LLMs to simulate how real patients react through the interplay of beliefs, thoughts, emotions, situations and other factors shown in Figure 2c. With high-fidelity simulations of patients, the ability to understand patients is trained: after interviewing simulated patients, novice therapists are required to independently describe the patient’s underlying cognitive model and receive feedback by comparing it to the standard answer—the cognitive model originally used to simulate their interviewee. This innovative yet safer training paradigm inspires us to explore a more sophisticated and practice-grounded training task aimed at improving the interview behaviors of therapists, specifically their ability to demonstrate effective, adaptive, and appropriate interactive behaviors during interviews, rather than focusing solely on post-session understanding of patients.
@@
However, a fundamental contradiction makes the training of therapeutic behaviors particularly challenging: as a training system, more “clear answers” as feedback are required to ensure consistent and goal-oriented learning within a low-risk and controlled environment (Arango-Ibanez et al. 2024)—much like the use of gold-standard cognitive models as feedback in the patient understanding training task. Yet, there is no absolute “gold standard” for appropriate interviewing behaviors in actual therapeutic practice (Leichsenring and Steinert 2017). The diversity and complexity of therapeutic approaches make it very difficult to provide unified, authoritative standards. Imposing a universal standard could lead to homogenized styles, suppress the individuality and creativity of therapists, and undermine their self-efficacy (Larson and Daniels 1998; Leichsenring and Steinert 2017). Conversely, certain mistaken therapy behaviors tend to be universal and relatively easy to identify (Gerke et al. 2020) such as neglect of the patient’s emotional expression, breaches of ethical standards, and transference or countertransference issues (Vyb´ıral et al. 2024). If we can leverage these relatively clear mistaken behaviors as labels to construct the training environment, then when trainees exhibit these behaviors, the corresponding targeted corrective principles can be retrieved and combined with the current context as relatively definite feedback to help them improve—our main idea to move a small step toward such open-ended training tasks.
@@
Existing work has begun to explore pipelines for creating realistic counselor-client dialogues for realistic counselor simulation. For example, the ESConv (Liu et al. 2021) dataset provides multi-turn emotional support dialogues to train supporter models using various strategies, while Chaszczewicz et al. (2024) augments a subset of ESConv with multi-layer feedback, linking strategy execution to evaluative feedback. Qiu and Lan (2024) employ role-play approaches to train LLMs for emotional support. Other efforts in more clinical scenarios, such as Li et al. (2024) and Shen et al. (2024) employ self-play or retrieval-based strategy to train LLM for the treatment of mental health disorders. However, these foundational client-counselor datasets are designed primarily to simulate supporters in realistic settings, rather than originally intended for training settings. As a result, the systematic customization and incorporation of training objectives and standards are often absent from the initial stages of dataset design.
@@
To train therapists in patient cognitive conceptualization, Patient-$\Psi$-Trainer (Wang et al. 2024) constructed patient cognitive models based on CBT principles, which serve as feedback standards for subsequent training. In this framework, LLMs simulate therapy patients, and after each interview, novice therapists are required to write out the patient’s underlying cognitive model, which is then compared to the standard answer for feedback. We adopt this innovative yet safer training paradigm; however, we explore a therapist training task with less standard answers, focusing on therapists’ generative abilities during the interview rather than merely assessing their understanding after the session.
@@
The framework is driven by two independent sets: the Mistaken Behavior Guideline $\mathbb{M}$ and the Client Case Set (Wang et al. 2024) $\mathbb{C}=\{C_{1},C_{2},...,C_{106}\}$. For the mistaken behavior therapist $T_{k}$, we instantiate a specialized therapist agent powered by carefully designed prompts for each $M_{k}\in\mathbb{M}$ (Figure 2b). This agent is configured to consistently exhibit the $k$-th specific mistake during conversations. Its behavior in the $i$-th dialogue turn is modeled as:
@@
Validate Stage: Compliance Verification. The refined feedback $f^{\prime}_{j,k}$ generated in the Refine stage is immediately audited by the same validator model $M_{V}$. Employing an “LLM-as-a-judge” framework (Lan et al. 2024), we establish validation criteria across four dimensions(progressivity, actionability, ethicality, supportiveness). The validator takes the dialogue $D_{j,k}$, the original feedback $f_{j,k}$, and the refined feedback $f^{\prime}_{j,k}$ as input, outputting a compliance judgment. The output is considered compliant only when all checklist criteria are met. Otherwise, the algorithm will re-attempt the refinement, up to a maximum of $N_{\text{retry}}$ times. If all attempts fail, the sample is flagged for subsequent human review.
@@
Legend: W. = Weighted; M. = Mean; EMR = Exact Match Ratio; $\Delta$ = Improvement of “(w/ Mate)” over the baseline.
@@
We performed parameter-efficient fine-tuning using LoRA (Low-Rank Adaptation) method on NVIDIA H800 GPU. The specific hyperparameter settings were as follows: $r=16$, LoRA scaling factor $\alpha=32$, LoRA dropout rate of $0.05$, the AdamW optimizer, learning rate of 0.0002, batch size of 4, gradient accumulation of 16 and epoch of 2.
@@
With Mate-tuned models achieving 30-40 BLEU and 25+ ROUGE-L improvements (Table 3), we examine whether gains translate to better supervisory content. Following the ”LLM-as-a-Judge” paradigm (Lee et al. 2024b), we use Deepseek-R1-0528 (Guo et al. 2025) for pairwise comparisons on 100 test samples across five criteria: Objectivity & Fairness, Constructiveness & Actionability, Professional Depth, Comprehensiveness, and Clarity & Structure.
@@
To investigate whether training on critical feedback tasks enhances broader psychological competencies, we evaluate models on empathy classification. Both feedback generation and empathy recognition require understanding emotional dynamics and appropriate therapeutic responses, making empathy detection a natural testbed for our therapeutically-trained models. We use the Empathy-Mental-Health (Sharma et al. 2020) dataset (911 samples) to classify whether responses demonstrate empathy in a zero-shot setting.
@@
To examine whether automated supervision strengthens novice counsellors’ self-efficacy(Larson and Daniels 1998), we built an online platform where participants conducted 15-minute text-based sessions with an LLM-driven virtual client. After each dialogue, a fine-tuned Supervisor model produced concrete improvement suggestions.
Six counselling-psychology graduate students (3 women, 3 men; average age $M_{\text{age}}=23$; all with fewer than ten hours clinical experience) completed the CASES-R(Hahn et al. 2021) twice: once before reading feedback and again after a five-minute reflection period. We focused on eight items from the exploration–insight and action subscales, each rated from 1 (no confidence) to 6 (complete confidence), yielding 48 paired observations. Scores increased significantly on all eight skills after feedback (10,000 sample bootstrap, 95% CIs). Gains were largest for procedural skills of Direct Guidance and Goal Setting, whereas Focus, already near the ceiling, improved only slightly. Interview data indicated that initial overconfidence in Reflection produced a ceiling effect that dampened its measurable improvement.

### v2

@@ Introduction
The field of mental health is grappling with a profound imbalance between the increasing number of patients and the scarcity of available therapists (WHO2023mentalhealth). With the rapid expansion of web-based mental health platforms and online counseling services, ensuring quality supervision and training in distributed, internet-mediated therapeutic environments has become a critical challenge for the web community.
Although Large language models (LLMs) offer a potential solution to this imbalance by simulating human therapist behaviors and providing personalized counseling services at lower costs (xiao2024healme; na2024cbt; xie2024psydt; wang2025emotional) , their direct application in patient-facing treatment poses several risks. These risks include ethical dilemmas (ong2024ethical), safety concerns (hager2024evaluation), potential harm with uncertainty about responsibility (nichol2023not), and the inherent challenge of establishing a therapeutic alliance with virtual therapists (choudhury2024large). The creation of structured, semantically-rich datasets that can support intelligent supervision systems represents a fundamental step toward scalable quality assurance in web-based therapeutic services. This prompts us to explore automated methods for training or supervising therapists to accelerate their development, addressing the imbalance by empowering more therapists through AI-assisted means.
@@
With significant advances in LLM role play  (xu2022cosplay; tseng2024two; chen2024persona), a new AI-assisted training paradigm has emerged (yang2024social). In this paradigm, LLMs play partners in the field and collaborate with practitioners to create interactive scenarios for experiential learning (kolb2014experiential). For example, in educational scenarios, “AI Student as Partner” helps teaching assistants improve their communication and classroom management skills (markel2023gpteach); in conflict resolution, “AI Opponent as Partner” helps users practice negotiation and conflict resolution (shaikh2023rehearsal). However, in psychotherapy, “AI Patient as Partner” (10948541) is difficult to implement. Privacy concerns and limited clinical data hinder the accurate modeling of patients’ behavioral patterns, as well as the already intricate cognitive states (reger2021virtual; lee-etal-2024-cactus).
@@
Recently, a pioneering work, Patient-$\Psi$-Trainer has been proposed to fill this gap (wang2024patient). It integrates 106 cognitive models of patients (beck2020cognitive), manually created by clinical experts, into LLMs to simulate how real patients react through the interplay of beliefs, thoughts, emotions, situations and other factors shown in Figure 2c. With high-fidelity simulations of patients, the ability to understand patients is trained: after interviewing simulated patients, novice therapists are required to independently describe the patient’s underlying cognitive model and receive feedback by comparing it to the standard answer—the cognitive model originally used to simulate their interviewee. This innovative yet safer training paradigm inspires us to explore a more sophisticated and practice-grounded training task aimed at improving the interview behaviors of therapists, specifically their ability to demonstrate effective, adaptive, and appropriate interactive behaviors during interviews, rather than focusing solely on post-session understanding of patients.
@@
However, a fundamental contradiction makes the training of therapeutic behaviors particularly challenging: as a training system, more “clear answers” as feedback are required to ensure consistent and goal-oriented learning within a low-risk and controlled environment (arango2024evidence)—much like the use of gold-standard cognitive models as feedback in the patient understanding training task. Yet, there is no absolute “gold standard” for appropriate interviewing behaviors in actual therapeutic practice (leichsenring2017cognitive). The diversity and complexity of therapeutic approaches make it very difficult to provide unified, authoritative standards. Imposing a universal standard could lead to homogenized styles, suppress the individuality and creativity of therapists, and undermine their self-efficacy (larson1998review; leichsenring2017cognitive). Conversely, certain mistaken therapy behaviors tend to be universal and relatively easy to identify (gerke2020frequencies) such as neglect of the patient’s emotional expression, breaches of ethical standards, and transference or countertransference issues (vybiral2024negative). If we can leverage these relatively clear mistaken behaviors as labels to construct the training environment, then when trainees exhibit these behaviors, the corresponding targeted corrective principles can be retrieved and combined with the current context as relatively definite feedback to help them improve—our main idea to move a small step toward such open-ended training tasks.
@@
Existing work has begun to explore pipelines for creating realistic counselor-client dialogues for realistic counselor simulation. For example, the ESConv (liu2021towards) dataset provides multi-turn emotional support dialogues to train supporter models using various strategies, while chaszczewicz2024multi augments a subset of ESConv with multi-layer feedback, linking strategy execution to evaluative feedback. qiu2024interactive employ role-play approaches to train LLMs for emotional support. Other efforts in more clinical scenarios, such as li2024mentalarena and shen2024large employ self-play or retrieval-based strategy to train LLM for the treatment of mental health disorders. However, these foundational client-counselor datasets are designed primarily to simulate supporters in realistic settings, rather than originally intended for training settings. As a result, the systematic customization and incorporation of training objectives and standards are often absent from the initial stages of dataset design.
@@
To train therapists in patient cognitive conceptualization, Patient-$\Psi$-Trainer (wang2024patient) constructed patient cognitive models based on CBT principles, which serve as feedback standards for subsequent training. In this framework, LLMs simulate therapy patients, and after each interview, novice therapists are required to write out the patient’s underlying cognitive model, which is then compared to the standard answer for feedback. We adopt this innovative yet safer training paradigm; however, we explore a therapist training task with less standard answers, focusing on therapists’ generative abilities during the interview rather than merely assessing their understanding after the session.
@@
The framework is driven by two independent sets: the Mistaken Behavior Guideline $\mathbb{M}$ and the Client Case Set (wang2024patient) $\mathbb{C}=\{C_{1},C_{2},...,C_{106}\}$. For the mistaken behavior therapist $T_{k}$, we instantiate a specialized therapist agent powered by carefully designed prompts for each $M_{k}\in\mathbb{M}$ (Figure 2b). This agent is configured to consistently exhibit the $k$-th specific mistake during conversations. Its behavior in the $i$-th dialogue turn is modeled as:
@@
Validate Stage: Compliance Verification. The refined feedback $f^{\prime}_{j,k}$ generated in the Refine stage is immediately audited by the same validator model $M_{V}$. Employing an “LLM-as-a-judge” framework (lan2024criticeval), we establish validation criteria across four dimensions(progressivity, actionability, ethicality, supportiveness). The validator takes the dialogue $D_{j,k}$, the original feedback $f_{j,k}$, and the refined feedback $f^{\prime}_{j,k}$ as input, outputting a compliance judgment. The output is considered compliant only when all checklist criteria are met. Otherwise, the algorithm will re-attempt the refinement, up to a maximum of $N_{\text{retry}}$ times. If all attempts fail, the sample is flagged for subsequent human review.
@@
Legend: W. = Weighted; M. = Mean; EMR = Exact Match Ratio; $\Delta$ = Improvement of “(w/ Mate)” over the baseline. Green indicates improvement, red indicates decline.
@@
We performed parameter-efficient fine-tuning using LoRA (Low-Rank Adaptation) method on NVIDIA H800 GPU. The specific hyperparameter settings were as follows: $r=16$, LoRA scaling factor $\alpha=32$, LoRA dropout rate of $0.05$, the AdamW optimizer, learning rate of 0.0002, batch size of 4, gradient accumulation of 16 and epoch of 2. Our experiments were conducted on open-source models with relatively small parameter counts, including Llama3.1-8B (grattafiori2024llama3herdmodels), Qwen3-8B/14B (yang2025qwen3). For baseline comparisons, we evaluated on the open-source Qwen2.5-1,5B/7B (yang2024qwen2), DeepSeek-V3-685B (guo2025deepseek) models and closed-source models GPT-4o and Claude-Sonnet-4. During inference, we set the maximum retry attempts $N_{\text{retry}}=3$, with temperature parameters configured as follows: client agent $T=0.15$, therapist agent $T=0.1$, and supervisor agent $T=0$.
To demonstrate the practical utility and validate the effectiveness of the Mate dataset, we developed a web-based feedback generation system for therapist training and supervision, offering a safe and controlled practice environment. As shown in Figure 3, the system comprises three core modules: simulated clients, a dialogue interface, and a supervisor module.
Figure 3: Therapist Training System. A web platform built on Mate with: (1) simulated clients, (2) a dialogue interface, and (3) an LLM-as-Supervisor that detects and localizes mistaken therapeutic behaviors (e.g., forcing change) and provides targeted feedback.
@@
With Mate-tuned models achieving 30-40 BLEU and 25+ ROUGE-L improvements (Table 3), we examine whether gains translate to better supervisory content. Following the ”LLM-as-a-Judge” paradigm (lee2024cactus), we use Deepseek-R1-0528 (guo2025deepseek) for pairwise comparisons on 100 test samples across five criteria: Objectivity & Fairness, Constructiveness & Actionability, Professional Depth, Comprehensiveness, and Clarity & Structure.
@@
To investigate whether training on critical feedback tasks enhances broader psychological competencies, we evaluate models on empathy classification. Both feedback generation and empathy recognition require understanding emotional dynamics and appropriate therapeutic responses, making empathy detection a natural testbed for our therapeutically-trained models. We use the Empathy-Mental-Health (sharma2020computational) dataset (911 samples) to classify whether responses demonstrate empathy in a zero-shot setting.
@@
To examine whether automated supervision strengthens novice counsellors’ self-efficacy(larson1998review; wang2025exploring; tang2025enhancing), we built an online web platform where participants conducted text-based sessions with an LLM-driven virtual client. After each dialogue, a fine-tuned supervisor model produced concrete improvement suggestions. As shown in Figure 3, a family relationship training case, the simulated client expresses distress about tension with her mother-in-law: “I feel like she’s always criticizing me… I don’t know what to do.” The therapist responds with advice: “Enduring will only let this situation continue, expressing your inner feelings would be better”.
Six counselling-psychology graduate students (3 women, 3 men; average age $M_{\text{age}}=23$; all with fewer than ten hours clinical experience) completed the CASES-R(hahn2021assessment) twice: once before reading feedback and again after a five-minute reflection period. We focused on eight items from the exploration–insight and action subscales, each rated from 1 (no confidence) to 6 (complete confidence), yielding 48 paired observations. Scores increased significantly on all eight skills after feedback (10,000 sample bootstrap, 95% CIs). Gains were largest for procedural skills of Direct Guidance and Goal Setting, whereas Focus, already near the ceiling, improved only slightly. Interview data indicated that initial overconfidence in Reflection produced a ceiling effect that dampened its measurable improvement.